What is the theoretical yield of PI3 from the reaction of 27.0 g of P and 68.0 g of I2?

2P(s) + 3I2(s) → 2PI3(s)


1
Expert's answer
2022-02-17T12:45:02-0500

Molar mass of I2 = 126.9 × 2 g/mol = 253.8 g/mol

Mol of I2 in 68.0 g = 68.0g/253.8 g/mol = 0.2679mol

From the balanced equation:

3 mol of I2 produce 2 mol of PI3

Then, 0.2679 mol of I2 produce = 0.2679 × 2/3 = 0.1786 mol PI3

Molar mass of PI3 = 411.69 g/mol

Mass of PI3 produced = 0.1786 mol × 411.69 g/mol = 73.53g of PI3

68.0 g of I2 produce 73.53g of PI3 i.e. theoretical yield of the reaction.